A hospitalized client is receiving continous iv fluids 1 Liter of Normal saline over 12 hrs. The client is also receiving a g - tube feeding at 45 ml/ hr with 59 ml flushes at 1000 and 1400. How much fluid intake ( ml) would the cleint receive for 8 hrs 0700-1800 day shift.
Answer #1

Given

Total volume in12 hours=1000ml

So in hour 1000÷12=83.3 ml

The start time for IV fluid is at 7am and ends at 6pm(1800)

For 11 hours (8×83.3=666.4)

So 666.4 ml of IV fluid is administered from 7am to 3pm (8hours)

Through g tube in the mean time

10am=59ml

2pm=59ml

Total g feed from 7 am to 3pm is 59ml+59ml=118

To calculate the total intake add IV intake with g -tube fee

=666.4 ml+118ml

=784.4 ml

The client will receive 784.4ml of fluid from 7am to 4pm


Time IVF G-tube
7am 83.3ml
8am 83.3ml
9am 83.3ml
10am 83.3ml 59ml
11am 83.3ml
12noon 83.3ml
1pm 83.3ml
2pm 83.3ml 59ml
3pm 83.3ml
Total 666.4 118ml

Total:IV+G tube

=666.4+118

=784.4ml